# BSRBF-KAN: A combination of B-splines and Radial Basis Functions in Kolmogorov-Arnold Networks

### v2

@@ Abstract
In this paper, we introduce BSRBF-KAN, a Kolmogorov Arnold Network (KAN) that combines B-splines and radial basis functions (RBFs) to fit input vectors in data training. We perform experiments with BSRBF-KAN, MLP, and other popular KANs, including EfficientKAN, FastKAN, FasterKAN, and GottliebKAN over the MNIST and Fashion-MNIST datasets. BSRBF-KAN shows stability in 5 training sessions with a competitive average accuracy of 97.55% on MNIST and 89.33% on Fashion-MNIST and obtains convergence better than other networks. We expect BSRBF-KAN to open many combinations of mathematical functions to design KANs. Our repo is publicly available at: https://github.com/hoangthangta/BSRBF-KAN.
@@ 1 Introduction
A recent work of  Liu et al. [1] over KANs opened a new paradigm in applying learnable activation functions as "edges" to fit training data instead of using fixed ones as "nodes" that are usually used in MLPs. The theory behind KANs relies on the Kolmogorov-Arnold representation theorem (KART), which states that a continuous function of multiple variables can be expressed as a combination of continuous functions of a single variable through additions. KANs are anticipated to bring a fresh perspective to solving issues that have been overshadowed by MLPs.
With that inspiration, many scientists have flocked to develop different types of KANs based on popular polynomial and basis functions. While these works focus on univariate functions to set up KANs, none have explored their combination. Therefore, we aim to combine B-splines [2] and Radial Basis Functions [3] to build a combined KAN named BSRBF-KAN. The reason why we use them is their successful implementation in two popular KANs, EfficientKAN [^1] https://github.com/Blealtan/efficient-kan and FastKAN [4]. Additionally, they are continuous basis functions used for function approximation, providing features such as smoothness, adaptability, and local control. With the experiments on MNIST [5] and Fashion-MNIST [6], we show the stability and faster convergence of BSRBF-KAN compared to other networks.
@@ 2 Related Works
In 1957, Kolmogorov provided proof to solve Hilbert’s 13th problem, demonstrating that a multivariate continuous function can be represented as a combination of single-variable functions and additions, known as KART [7, 8]. This theorem has been applied in numerous works to build neural networks [9, 10, 11, 12, 13, 14]. However, it has also sparked debates about its ability in designing neural networks [15, 16].
Despite the long history of application in neural networks, no works on KART had gained significant attention within the research community until the recent work by Liu et al. [1]. They recommended not adhering strictly to the KART but generalizing it to create KANs with more neurons and layers. As a result, their KANs can surpass MLPs in both accuracy and interpretability on small-scale AI + Science tasks.
@@ 3.2 The design of KANs
Liu et al. [1] was inspired by KART to design KANs, but authors advised generalizing it to more widths and depths. In Equation 1, we must search proper $\Phi_{q}$ and $\phi_{q,p}$ to solve the problem. A general KAN network consisting of $L$ layers takes $\mathbf{x}$ to generate the output as:
@@ 3.3 Implementation of the current KANs
FasterKAN outperforms FastKAN in both forward and backward processing speeds [17]. It uses Reflectional Switch Activation Functions (RSWAFs), which are straightforward to compute due to their uniform grid structure. The RSWAF function $\phi(r)$ is shown as:
@@
GottliebKAN showed the best result compared to the other 17 polynomial basic functions for setting up KANs on MNIST [21]. Gottlieb polynomials are a family of polynomials that arise in the study of the Bernoulli numbers and have applications in combinatorics and number theory [21].
@@ 3.4 BSRBF-KAN
where $b(x)$ and $w_{b}$ are the base output and its base matrix. $\phi_{BS}(x)$ and $\phi_{RBF}$ are B-Spline and RBF functions, and $w_{s}$ is the spline matrix. The design of BSRBF-KAN is similar to the original KAN, mentioned in Section 3.2. Given the goal of designing a combined KAN, our intuition understands that the forward and backward speeds will not surpass those of individual KANs.
@@ 4 Experiments
We conducted 5 training sessions for each model on MNIST [5] and Fashion-MNIST [6] to assess their overall performance more reliably. By taking the best metric value from all sessions, we aimed to mitigate the effects of variability in training and ensure that the evaluation reflects the models’ peak capabilities. With the intention of designing simple networks, we only use activation functions, linear transformation, and layer normalization.
All models have a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). For GottliebKAN, we used a structure (784, 64, 64, 10) to make it the same as the original design [^5] https://github.com/seydi1370/Basis_Functions. We trained the models with 15 epochs on MNIST and 25 epochs on Fashion-MNIST. Other hyperparameters are the same in all models, including batch_size = 64, learning_rate = 1e-3, weight_decay=1e-4, optimize = AdamW, and loss=CrossEntropy.
Figure 1 demonstrates training losses of models represented by logarithm values in a certain training time. On both MNIST and Fashion-MNIST, the training loss values of BSRBF-KAN drop sharply continuously, which makes it the model with the smallest loss after the training process. In contrast, GottliebKAN lost the least loss, which affected its performance in the next experiments.
Figure 1: The logarithm values of training losses (left) and validation losses (right) in 15 epochs in a training session.
Table 1: The best metric values in 5 training sessions on MNIST and Fashion-MNIST.
@@
Table 2: The average metric values in 5 training sessions on MNIST and Fashion-MNIST.
@@ 4.4 Ablation Study
Several ablation experiments were performed to help us understand the impact of various components within BSRBF-KAN on overall model performance. Table 3 presents the metric values for different combinations of BSRBF-KAN components. Initially, we trained the model with all components included, followed by training variants with some components removed. The results show that the base output and layer normalization are key for the model’s performance, as their both absence leads to the worst results.
@@
Table 3: The performance of BSRBF-KAN by different components on MNIST and Fashion-MNIST in a training setup.
@@ 5 Conclusion
We introduced BSRBF-KAN, a novel KAN that integrates B-splines and Radial Basis Functions (RBFs) to fit input vectors in the training process. We did not intend to make BSRBF-KAN better than MLP and other KANs by optimizing hyperparameters; instead, we built it by default to understand aspects of KANs combination in training models. In the experiments, we trained BSRBF-KAN, MLP, and other KANs on the MNIST and Fashion-MNIST datasets. BSRBF-KAN demonstrated competitive performance and rapid convergence compared to other models. Furthermore, it is the best model if counting on the average values of both datasets. However, its high convergence rate can lead to overfitting, which can be mitigated by using dropout.

### v3

@@ Abstract
In this paper, we introduce BSRBF-KAN, a Kolmogorov Arnold Network (KAN) that combines B-splines and radial basis functions (RBFs) to fit input vectors during data training. We perform experiments with BSRBF-KAN, multi-layer perception (MLP), and other popular KANs, including EfficientKAN, FastKAN, FasterKAN, and GottliebKAN over the MNIST and Fashion-MNIST datasets. BSRBF-KAN shows stability in 5 training runs with a competitive average accuracy of 97.55% on MNIST and 89.33% on Fashion-MNIST and obtains convergence better than other networks. We expect BSRBF-KAN to open many combinations of mathematical functions to design KANs. Our repo is publicly available at: https://github.com/hoangthangta/BSRBF_KAN.
@@ 1 Introduction
A recent work of  Liu et al. [1] mentioned applying learnable activation functions as "edges" to fit training data instead of using fixed ones as "nodes" that are usually used in multi-layer perceptions (MLPs). The theory behind KANs relies on the Kolmogorov-Arnold representation theorem (KART), which states that a continuous function of multiple variables can be expressed as a combination of continuous functions of a single variable through additions [2]. KANs are anticipated to bring a fresh perspective to solving issues that have been overshadowed by MLPs.
With that inspiration, many scientists have flocked to develop different types of KANs based on popular polynomials and basis functions. While these works focus on univariate functions to set up KANs, none have explored their combination. Therefore, we aim to combine B-splines [3] and RBFs [4] to build a combined KAN named BSRBF-KAN. The reason why we use them is their successful implementation in two popular KANs, EfficientKAN [^1] https://github.com/Blealtan/efficient-kan and FastKAN [5]. Additionally, they are continuous basis functions used for function approximation, providing features such as smoothness, adaptability, and local control. With the experiments on MNIST [6] and Fashion-MNIST [7], we show the stability and faster convergence of BSRBF-KAN compared to other networks. In this paper, our primary contribution is the development of a combined KAN, BSRBF-KAN, which we compare with other networks on two popular datasets (MNIST and Fashion-MNIST) to understand our KAN behaviors in the image classification problem.
The rest of this document is structured as follows: Section 2 contains related works on KART and KANs. Section 3 outlines our methodology, describing the KART and the design of the KAN architecture, as well as several currently popular KANs. Section 4 presents our experiments, where we compare BSRBF-KAN with MLP and other KANs using results from the MNIST and Fashion-MNIST datasets. This section also includes ablation studies to identify the key components of BSRBF-KAN. Finally, Section 5 provides conclusions, comments, and our future work.
@@ 2 Related Works
In 1957, Kolmogorov provided proof to solve Hilbert’s 13th problem, demonstrating that a multivariate continuous function can be represented as a combination of single-variable functions and additions, known as KART [2, 8]. This theorem has been applied in numerous works to build neural networks [9, 10, 11, 12, 13, 14]. However, it has also sparked debates about KART’s ability in designing neural networks. Girosi and Poggio [15] indicated that KART is irrelevant to neural networks because the inner function $\phi_{q,p}$ in Equation 1 can be highly non-smooth [16], preventing $f$ from being smooth, which is essential for generalization and noise stability in neural networks. In contrast, Kůrková [17] argued that KART is relevant to neural networks, demonstrating that all one-variable functions can be accurately approximated by linear combinations of affine functions with some sigmoidal functions. This suggests that while KART may sometimes be unsuitable for neural networks due to its potential non-smooth functions, there are cases where it can be effectively used if appropriate functions are identified.
Despite the long history of application in neural networks, no works on KART had gained significant attention within the research community until the recent work by Liu et al. [1]. They recommended not adhering strictly to the KART but generalizing it to create KANs with more neurons and layers. Our intuition thinks this is an important point to avoid the non-smooth problem when applying KART to neural networks. As a result, KANs can surpass MLPs in both accuracy and interpretability on small-scale AI + Science tasks. However, the revival of KAN has also faced criticism by Dhiman [18], where KANs are described as MLPs with spline-based activation functions, unlike traditional MLPs, which have fixed activation functions. This may offer insights that enhance our understanding of KANs and their comparison with MLPs.
@@ 3.2 The design of KANs
Let’s revisit an MLP characterized by affine transformations and non-linear functions. From an input $\mathbf{x}$, the network performs the composition of weight matrices by layers (from layer 00 to layer $L-1$) and the non-linearity $\sigma$ to generate the final output.
$\text{MLP}(\mathbf{x})$ $=(W_{L-1}\circ\sigma\circ W_{L-2}\circ\sigma\circ\cdots\circ W_{1}\circ\sigma\circ W_{0})\mathbf{x}$ (2)
$=\sigma\left(W_{L-1}\sigma\left(W_{L-2}\sigma\left(\cdots\sigma\left(W_{1}\sigma\left(W_{0}\mathbf{x}\right)\right)\right)\right)\right)$
Liu et al. [1] was inspired by KART to design KANs, but the authors advised generalizing it to more widths and depths. In Equation 1, we must search proper $\Phi_{q}$ and $\phi_{q,p}$ to solve the problem. A general KAN network consisting of $L$ layers takes $\mathbf{x}$ to generate the output as:
@@ 3.3 Implementation of the current KANs
FasterKAN outperforms FastKAN in both forward and backward processing speeds [19]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function $\phi(r)$ is shown as:
@@
GottliebKAN, which is based on Gottlieb polynomials, showed the best result compared to the other 17 polynomial basic functions for setting up KANs on MNIST [23]. Gottlieb polynomials are a family of polynomials that arise in the study of the Bernoulli numbers and have applications in combinatorics and number theory [26]. Teymoor Seydi [23] denoted Gottlieb polynomials as:
$G_{n}(\mathbf{x})=\sum_{k=0}^{n}\binom{n}{k}B_{k}(\mathbf{x})\mathbf{x}^{n-k}$ (12)
where $G_{n}(\mathbf{x})$ represents the $n$-th Gottlieb polynomial, $(n\;k)$ denotes the binomial coefficient "n choose k", $B_{k}(\mathbf{x})$ are the Bernstein basis polynomials. Finally, $\mathbf{x}^{n-k}$ is the power term that adjusts the degree of $B_{k}(\mathbf{x})$ within the polynomial $G_{n}(\mathbf{x})$.
@@ 3.4 BSRBF-KAN
where $b(x)$ and $w_{b}$ are the base output and its base matrix. $\phi_{BS}$ and $\phi_{RBF}$ are B-Spline and RBF functions, and $w_{s}$ is the spline matrix. Given the goal of designing a combined KAN, our intuition understands that the forward and backward speeds will not surpass those of individual KANs.
@@ 4 Experiments
Figure 1: The logarithm values of training losses (left) and validation losses (right) in 15 epochs in a training run.
We conducted 5 training runs for each model on MNIST [6] and Fashion-MNIST [7] to assess their overall performance more reliably. By taking the best metric value from all runs, we aimed to mitigate the effects of variability in training and ensure that the evaluation reflects the models’ peak capabilities. With the intention of designing simple networks, we only use activation functions, linear transformation, and layer normalization.
All models have a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). For GottliebKAN, we used a structure (784, 64, 64, 10) to make it the same as the original design [^5] https://github.com/seydi1370/Basis_Functions. We trained the models with 15 epochs on MNIST and 25 epochs on Fashion-MNIST. Other hyperparameters are the same in all models, including batch_size=64, learning_rate=1e-3, weight_decay=1e-4, gamma=0.8, optimize=AdamW, and loss=CrossEntropy.
Figure 1 demonstrates training losses of models represented by logarithm values in a certain training run. On both MNIST and Fashion-MNIST, the training loss values of BSRBF-KAN drop sharply continuously, which makes it the model with the smallest loss after the training process. While this accelerates the training, there is also a concern that the model might be overfitting. We examined both training and validation accuracies across datasets and found no anomalies. In contrast, GottliebKAN lost the least loss, which affected its performance, as shown in the next experiments.
Table 1: The best metric values in 5 training runs on MNIST and Fashion-MNIST.
@@
Table 2: The average metric values in 5 training runs on MNIST and Fashion-MNIST.
@@ 4.4 Ablation Study
Several ablation experiments were performed to help us understand the impact of various components within BSRBF-KAN on overall model performance. Table 3 presents the metric values for different combinations of BSRBF-KAN components. Initially, we trained the model with all components included, followed by training variants with some components removed. The results indicate that the base output and layer normalization are key for the model’s performance, as their both absence leads to the worst results. The base output
@@
Table 3: The performance of BSRBF-KAN by different components on MNIST and Fashion-MNIST in a training run.
@@ 5 Conclusion
We introduced BSRBF-KAN, a novel KAN that integrates B-splines and RBFs to fit input data in the training process. We did not intend to make BSRBF-KAN better than MLP and other KANs by optimizing hyperparameters; instead, we built it by default to understand aspects of KANs combination in training models. In the experiments, we trained BSRBF-KAN, MLP, and other KANs on the MNIST and Fashion-MNIST datasets. BSRBF-KAN demonstrated competitive performance and rapid convergence compared to other models. Furthermore, it is the best model if counting on the average values of both datasets. However, its high convergence rate can lead to overfitting, which can be mitigated by using dropout and other techniques.